The demand function for rose is Q=a-bp, and the supply function is Q=c+ep+ft, where a, b, c, e, and f are positive constants and t is the average temperature in a month. Show how the equilibrium quantity and price vary with temperature?
1
Expert's answer
2015-06-04T00:00:45-0400
Qd = a­p-b, Qs = c+ep+ft. In equilibrium Qd = Qs, so: 
a­pb = c+ep+ft
p*(e + b) = a - c - ft
p = (a - c - ft)/(e + b)
So, equilibrium price increases, when the temperature falls and equilibrium quantity falls too.